In the following problem, check that it is appropriate to use the normal approximation to the binomial. Then use the normal distribution to estimate the requested probabilities What are the chances that a person who is murdered actually knew the murderer? The answer to this question explains why a lot of police detective work begins with relatives and friends of the victim! About 63% of people who are murdered actually knew the person who committed the murder. T Suppose that a detective file in New Orleans has 58 current unsolved murders. Find the following probabilities. (Round your answers to four decimal places.) (a) at least 35 of the victims knew their murderers (b) at most 48 of the victims knew their murderers (c) fewer than 30 victims did not know their murderers d) more than 20 victims did not know their murderers
Answer #1
  • Here X is the random sample denoting the number of victims who knew their murderer
  • Since here samples follow normal distribution we will use normal approximation
  • The estimated proportion is given by p^=x/n
  • The standard error is given by o) =0.63(1 - 0.63) 58=0.6339
  • To calculate the probability atleast 35 victim of them knew there murderer
  • P(Z>P-p l-p) 7t)
  • =P(Z>35/58 -0.63 0.63 (1-0.63))
  • =P(Z>-0.4188)=P(Z<0.4188)
  • =0.65910 (i.e the probability value associated with the z value in the normal table
  • b)
  • To calculate the probability atmost 48 victimeof them knew there murderer is
  • P(Z>P-p l-p) 7t)
  • =P(Z>48/58 0.63 0.63 (1-0.63))
  • =P(Z>3.1167)
  • =1-P(Z<3.1167)
  • =1-.99906
  • =0.00094
  • c)
  • To calculate the probability fewer than 30 victim knew there murderer
  • P(Z<P-p l-p) 7t)
  • P(Z<30/58 -0.63 0.63 (1-0.63))
  • P(Z<-1.7786)
  • .03836
  • To calculate the probability more than 20 victim knew there murderer
  • P(Z>P-p l-p) 7t)
  • P(Z>20/58 -0.63 0.63 (1-0.63))
  • =P(Z>-4.4983)
  • =P(Z<4.4983)
  • =.99997